OpenStudy (anonymous):

Write an equation in point slope form of the line that passes through (5, 0), and whose slope is -2. y = -2(x – 5) y – 5 = -2(x - 5) y = -2(x + 5) y = 2(x – 5)

OpenStudy (anonymous):

It will be in the form y=kx+c. What will be k in this case?

OpenStudy (anonymous):

no idea ill give you a medal for the answer i got sooo many questions left and just some of this stuff i dont get so plzz just answers sorry to be snappy but after asnwerin 150 you really dont want to do the other 50

OpenStudy (anonymous):

k=gradient=rise/run=-2 So y=-2+c what is c? insert (5,0): 0=-2(5)+c 0=-10+c c=10

OpenStudy (anonymous):

*y=-2x+c

OpenStudy (anonymous):

so is it the first answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

<3